# A Field Guide to Event-Shape Observables Using Optimal Transport

### v1

@@ Abstract
We lay out the phenomenological behavior of event-shape observables evaluated by solving optimal transport problems between collider events and reference geometries—which we name manifold distances—to provide guidance regarding their use in future studies. This discussion considers several choices related to the metric used to quantify these distances. We explore the differences between the various options, using a combination of analytical studies and simulated minimum-bias and multi-jet events. Making judicious choices when defining the metric and reference geometry can improve sensitivity to interesting signal features and reduce sensitivity to non-perturbative effects in QCD. The goal of this article is to provide a ‘field guide’ that can inform how choices made when defining a manifold distance can be tailored for the analysis at-hand.
@@ 1 Introduction
Event-shape observables are not universally discriminating; the choice of most effective observable is determined by the physics of interest to the analysis at-hand. For example, thrust PhysRevLett.39.1587 ; BRANDT196457 ; DERUJULA1978387 is often the preferred observable for analyses of QCD at high momentum transfers. Thrust is a motivated choice in this case, as it quantifies how similar an event is to a $2\rightarrow 2$ point-like particle scattering geometry. However, for physics scenarios with non-jetty topologies (such as homogeneous/uniform radiation patterns, high jet multiplicities, etc., thrust can be less effective for selecting signal events than other options. We see these kinds of radiation patterns are observed in the radiation of black holes produced at colliders Giddings:2001bu ; Dimopoulos:2001hw ; Meade:2007sz , multi-jet final states in supersymmetric scenarios Evans:2013jna ; Barbier:2004ez , high-multiplicity and strongly coupled new physics scenarios Strassler:2006im ; Polchinski:2002jw ; Hofman:2008ar ; Strassler:2008fv ; Strassler:2008bv ; Hatta:2008tx , and models that generically fall under the category of soft, unclustered energy patterns (SUEPs) Knapen:2016hky .
@@
A broad library of event shapes can facilitate investigations into diverse event geometries, helping to ensure that no stone is left unturned in the search for BSM physics at the LHC and at future colliders Narain:2022qud . In this work, we examine the phenomenology of event-shape observables that are defined using a recently-proposed methodology of quantifying the distance of collider events to symmetric reference geometries in terms of a formal metric. These observables are computed using the particle physics application of the Earth Mover’s Distance 10.1007/978-3-642-40020-9_43 ; 192468 ; Rubner:1998:MDA:938978.939133 ; Rubner2000 ; 10.1007/978-3-540-88690-7_37 – or Wassterstein distance wasserstein1969markov ; dobrushin1970prescribing – called the Energy Mover’s Distance (EMD) Komiske:2019fks ; Komiske:2020qhg . We will generically refer to these observables as manifold distances. The first application of the EMD in the context of designing new event shapes was the definition of event isotropy Cesarotti:2020hwb , where the reference geometry is a (quasi-) isotropic radiation pattern. Canonical event shapes like thrust, sphericity PhysRevD.1.1416 ; Ellis:1976uc and spherocity PhysRevLett.39.1237 were shown to be less efficient discriminants for signals with quasi-isotropic signatures Cesarotti:2020ngq . Event isotropy was recently measured in proton-proton collisions at the Large Hadron Collider in high-$p_{T}$ multijet events by the ATLAS Collaboration STDM-2020-20 , and in minimum-bias events by the CMS Collaboration CMS:2024bwe . Additionally the EMD has been used to identify jet-substructures Ba:2023hix ; Kitouni:2022qyr , and optimal transport more generally has been proposed to further study collider events and identify anomalies Cai:2020vzx ; Cai:2021hnn ; Craig:2024rlv .
@@
Here, we build on the foundation laid by event isotropy and address questions relevant for more generalized event shape definitions using EMDs. We do so by varying the two fundamental components that specify the observable: the reference geometry and the distance metric. [^1] One could also change the energy dependence, which is related to changing the power in computing in the sphericity tensor PhysRevD.1.1416 ; Ellis:1976uc or the $p_{T}^{D}$ observable CMS-HIG-11-027 . As this change would result in the loss of IRC safety, we have not explored this topic in this work. We consider two complementary geometries in the transverse plane: an $N=2$ back-to-back particle event $\mathcal{E}_{2}$ from the manifold of such events $\mathcal{M}_{2}$, and a discrete, uniform ‘ring-like’ radiation pattern $\mathcal{U}_{N}$. Additionally we consider several different distance metrics to illustrate how the choice of metric can affect the sensitivity.
@@ 2 Definition of event observables with EMD
This construction can be generalized to probe other global event geometries or specific structures by comparing the collider event to geometries reference geometries. While there are many interesting candidates, we focus on two extremal examples: isotropic and dipole-like (back-to-back), as illustrated in Fig. 1. In the original work, the reference geometry was chosen to be a uniform (or, quasi-uniform, for computational tractability) distribution Cesarotti:2020hwb , but we will also consider the dipole geometry that was first shown in Ref. STDM-2020-20 to be an observable comparable to thrust. Additionally, we consider several distance measures to demonstrate how the sensitivity of these observables can be tuned to identify specific features of a geometry, e.g. by only considering transverse information in hadron collisions due to the unknown boost along the beamline. [^3] For the interested reader: Ref. STDM-2020-20 conducts some studies with longitudinal information included.
@@ 2.1 Reference Geometries
where $N$ is sufficiently large to approximate the infinite particle distribution. For this work, we use $N=256$, such that $\mathcal{U}\approx\mathcal{U}_{256}$, shown in Fig. 2. We suppress these subscripts for notational simplicity, and will refer to the quasi-isotropic event as $\mathcal{U}$.
@@
where $\theta$ is the relative angle of the event, and the factors of $1/2$ are from normalizing each particle energy by the total energy. One additional aspect of determining the dipole distance $\mathcal{D}_{\mathcal{M}_{2}}$is to minimize the EMD over the orientation of the reference geometry with respect to the collider event. This is analogous—but not identical—to finding the thrust axis, or the axis along which the event is most collimated. In practice, the minimization over this angle is performed numerically using standard libraries when computing the manifold distance.
@@ 2.4 Analytics
Figure 5: Manifold distances as a function of $N$-particle symmetric events (meaning equal energy, evenly distributed particles). The values of event isotropy are shown on the left and exhibit a monotonic behavior, all maximized for the 2-particle ($\mathcal{M}_{2}$) configuration, decreasing as the $N$-particle vent approaches the continuous limit. The values of (unnormalized) dipole distance $\mathcal{D}_{\mathcal{M}_{2}}$, shown on the right, are not monotonic, and jump around before asymptoting to the infinite particle limit. We show the unnormalized dipole distance values simply for ease of viewing. The dashed lines indicate the value of dipole distance computed for $N=\infty$. To the right of the solid line we only plot the value for $N$ divisible by 2 for ease of analytic computation.
The range of the manifold distances is closed and well defined for events with sets of massless particles and zero net momentum. Therefore, for ease of interpretation, we normalize the observable to the maximal value such that the observable runs from 0 to 1. However, it is not immediately obvious what the maximally distant radiation pattern is. To show this, we plot the manifold distances as a function of symmetric $N$-particle radiation patterns $\mathcal{U}_{N}$ in Fig. 5
@@
The computation is less straightforward for dipole distance. For distance measures of $\beta\geq 1$, the maximally-distant configuration to the back-to-back reference geometry is the three-prong event. However, for $\beta=0.5$, we see in Fig. 5b that the maximally distance configuration from the dipole event is the uniform radiation pattern.
@@
¯$\bar{\mathcal{I}}\equiv 1-\mathcal{I}$ (11)
@@ 3.1 Monte Carlo setups and selections
The first sample consists of groups generated using the default Pythia group of hard QCD $2\rightarrow 2$ processes, which we call the ‘Hard QCD’ (‘HQCD’) sample. The implementation of the anti-$k_{t}$ algorithm with radius $R=0.4$ Cacciari:2008gp from FastJet is used Fastjet ; Cacciari:2005hq to reconstruct jets from stable final-state particles with a transverse momentum above $0.5$ MeV and within $|y|<4.9$, excluding neutrinos. The events in this sample are required to have at least two anti-$k_{t}$ jets with a $p_{T}$ above $60$ GeV and rapidity $|y|<4.5$ to be included in the study. Additional jets are also retained if they are present and satisfy the same criteria.
@@
In this figure we note a few important features of the distributions. Firstly, as expected, both observables demonstrate that the SQCD sample is more isotropic than the HQCD sample. Secondly, the width of the distribution is larger for the choice of reference geometry that more closely approximates the radiation pattern – e.g. the spread is larger for SQCD with event isotropy than dipole distance, and the reverse is true for HQCD. This illustrates how important the choice of reference geometry is for signal discrimination. As we have stated previously, manifold distances (and similar observables like thrust) have a greater dynamic range when the distance is computed with respect to a similar reference geometry than a very distant one: being ‘far’ from a radiation pattern is often a more degenerate mapping. Finally, note that the distributions of the ‘further geometry’ are pushed more towards the extremal values, as indicated by the dashed lines at the mean of the distribution in Fig. 6. This feature reinforces the previous point, that ‘further’ events tend to saturate more quickly than ‘nearby’ events.
@@
While we include $\beta=4$ for the sake of study, the fact that we already see saturation at perfectly isotropic for event isotropy in the soft QCD sample suggests that using large $\beta$s can identify if there is approximate underlying symmetries in the set of events.
@@
We start by looking at the HQCD samples in Fig. 8. It is immediately evident that for all values of $\beta$, the distributions are much more correlated for dipole distance than for event isotropy. We conjecture that this is due to the large underlying symmetry of the reference geometry for event isotropy. On average the distances traveled by particles to get to the reference geometry of a dipole rather than the isotropic distribution are greater, and the non-linearities in the EMD calculation are more evident.
@@
We also note the general trend that when the values of $\beta$ are more different, the 2d distribution is broader. This is a statement underscoring how different choices of $\beta$ are fundamentally measuring different features of the radiation pattern: high $\beta$ is insensitive to small movements, whereas smaller $\beta$ weights small and large distances increasingly equally. The widening of the spread demonstrates that, depending on the analysis at hand, a different choice of $\beta$ might be a more discriminating probe.
Figure 8: Plots of the 2d distributions of $\mathcal{D}_{\mathcal{M}_{2}}$ (top) and $\bar{\mathcal{I}}$ (bottom) for the HQCD sample. We also plot the dashed curve for the $\bar{\mathcal{I}}$  of $\mathcal{U}_{N}$ on the $\bar{\mathcal{I}}$  as a simple analytic model of the underlying behavior. We see that $\bar{\mathcal{I}}$ is always more correlated, and that the correlations are less between high $\beta$ and low $\beta$.
@@
Now let’s move our attention to the SQCD samples to confirm our statements are generic to the observable, and not specific to dijet-like radiation patterns. The same plots are shown for SQCD in Fig. 9. Again, we generally find that dipole distance is less correlated than event isotropy, except for the case of $\beta=0.5$ and $\beta=1$. The argument for this behavior is the same as that for the HQCD samples, but with the difference that for sufficiently small $\beta$, even little distances become relevant in the event isotropy calculation when the collider event is also somewhat isotropic. As we increase $\beta$, the distance weight for nearby particles becomes so small that the distributions become highly correlated. From the plots in Fig. 9, we see that the correlation collapses to the curve for the maximally isotropic value for a fixed number of particles.
Overall, we can conclude that the choice of $\beta$ when computing event isotropy doesn’t change the analysis as significantly as the choice of $\beta$ for dipole distance. The biggest exception we see to this statement is small $\beta$ for quasi-isotropic samples can produce different results as it is more sensitive to small perturbations from the reference geometry. Therefore, we recommend that unless there is a particular motivation to use a specific $\beta$, one should use whichever is least computationally expensive for event isotropy. However, the choice of $\beta$, especially large vs. small, can noticeably change the distribution of dipole distance. The choice of $\beta$ should then be explored to optimize the analysis at hand.
@@
Now let’s fix the value of $\beta$ and see how correlated event isotropy and dipole distance are for each physics sample. The results are shown in Fig. 10. We see that for all $\beta$s and QCD samples, the observables are never exactly correlated. Some amount of correlation is expected as both observables indicate the underlying event shape, but the value of one observable is not predictive of the other.
@@ 3.4 Particle and Jet Multiplicity
$f(\bar{\mathcal{I}},N)=\frac{\mathcal{I}}{N}$ (12)
that is just the ratio of the event isotropy $\bar{\mathcal{I}}$ to the number of particles $N$ is a highly efficient measure. As shown in Fig. 12, we see that the spectra of the HQCD and SQCD sample are nearly completely separated, with an AUC value $>$ 0.99. We tested similar measures for the ratio and product of $\bar{\mathcal{I}}$ and $\mathcal{D}_{\mathcal{M}_{2}}$, but improvements to the AUC were negligible. A more complicated observable (e.g. one constructed with machine learning techniques hepmllivingreview ) may still exist. Generally, when observables can probe different features of events, a new function with the observables as inputs of can serve as a highly efficient characterization measure Thaler:2010tr ; Moult:2016cvt ; Komiske:2017aww .
Figure 12: The HQCD and SQCD spectra of the ratio between event isotropy $\bar{\mathcal{I}}$ and particle number $N$.
Alternatively, jet multiplicity reflects a more global picture of the radiation pattern as jets have a finite size radius. Thus a high-jet event suggests that radiation has gone in several distinct directions, whereas a high-multiplicity event could indicate only that there are many collimated particles in a single jet. To study this correlation, we plot the distribution of the HQCD sample with $N_{j}\geq$ 2, 3, 5, and 7 in Fig. 13 along with the SQCD sample for reference. We see that jet multiplicity tracks the global features of the event more reliably, as the distribution in both event-shape observables shifts towards the SQCD sample, reinforcing one’s intuition that higher jet-multiplicity indicates a more isotropic event.

### v2

@@ Abstract
We lay out the phenomenological behavior of event-shape observables evaluated by solving optimal transport problems between collider events and reference geometries—which we name manifold distances—to provide guidance regarding their use in future studies. This discussion considers several choices related to the metric used to quantify these distances. We explore the differences between the various options, for the first time using a combination of analytical studies and simulated minimum-bias and multi-jet events. Making judicious choices when defining the metric and reference geometry can improve sensitivity to interesting signal features and reduce sensitivity to non-perturbative effects in QCD. The goal of this article is to provide a ‘field guide’ that can inform how choices made when defining a manifold distance can be tailored for the analysis at-hand.
@@ 1 Introduction
Event-shape observables are not universally discriminating; the choice of most effective observable is determined by the physics of interest to the analysis at-hand. For example, thrust PhysRevLett.39.1587; BRANDT196457; DERUJULA1978387 is often the preferred observable for analyses of QCD at high momentum transfers. Thrust is a motivated choice in this case, as it quantifies how similar an event is to a $2\rightarrow 2$ point-like particle scattering geometry. However, for physics scenarios with non-jetty topologies (such as homogeneous/uniform radiation patterns, high jet multiplicities, etc., thrust can be less effective for selecting signal events than other options. These kinds of radiation patterns are created by collider-produced black holes Giddings:2001bu; Dimopoulos:2001hw; Meade:2007sz, multi-jet final states in supersymmetric scenarios Evans:2013jna; Barbier:2004ez, high-multiplicity and strongly coupled new physics scenarios Strassler:2006im; Polchinski:2002jw; Hofman:2008ar; Strassler:2008fv; Strassler:2008bv; Hatta:2008tx, and models that generically fall under the category of soft, unclustered energy patterns (SUEPs) Knapen:2016hky.
@@
A broad library of event shapes can facilitate investigations into diverse event geometries, helping to ensure that no stone is left unturned in the search for BSM physics at the LHC and at future colliders Narain:2022qud. In this work, we examine the phenomenology of event-shape observables that are defined using a recently-proposed methodology of quantifying the distance of collider events to symmetric reference geometries in terms of a formal metric. These observables are computed using the particle physics application of the Earth Mover’s Distance 10.1007/978-3-642-40020-9_43; 192468; Rubner:1998:MDA:938978.939133; Rubner2000; 10.1007/978-3-540-88690-7_37 – or Wasserstein distance wasserstein1969markov; dobrushin1970prescribing – called the Energy Mover’s Distance (EMD) Komiske:2019fks; Komiske:2020qhg. We will generically refer to these observables as manifold distances. The first application of the EMD in the context of designing new event shapes was the definition of event isotropy Cesarotti:2020hwb, where the reference geometry is a (quasi-) isotropic radiation pattern. Canonical event shapes like thrust, sphericity PhysRevD.1.1416; Ellis:1976uc and spherocity PhysRevLett.39.1237 were shown to be less efficient discriminants for signals with quasi-isotropic signatures Cesarotti:2020ngq. Event isotropy was recently measured in proton-proton collisions at the Large Hadron Collider in high-$p_{T}$ multijet events by the ATLAS Collaboration STDM-2020-20, and in minimum-bias events by the CMS Collaboration CMS:2024bwe. Additionally the EMD has been used to study jet substructure by proposing novel observables Ba:2023hix; Kitouni:2022qyr, and optimal transport more generally has been proposed to further study collider events and identify anomalies in the context of unsupervised learning studies Cai:2020vzx; Cai:2021hnn and novel triggering strategies Craig:2024rlv [^1] Reviews of jet substructure and anomaly detection studies at the LHC may be found respectively in Refs. Kogler:2018hem and Kasieczka:2021xcg for readers unfamiliar with these subfields..
@@
Here, we build on the foundation laid by event isotropy and address questions relevant for more generalized event shape definitions using EMDs. We do so by varying the two fundamental components that specify the observable: the reference geometry and the distance metric. [^2] One could also change the energy dependence, which is related to changing the power in computing in the sphericity tensor PhysRevD.1.1416; Ellis:1976uc or the $p_{T}^{D}$ observable CMS-HIG-11-027. As this change would result in the loss of IRC safety, we have not explored this topic in this work. We consider two complementary geometries in the transverse plane: an $N=2$ back-to-back particle event $\mathcal{E}_{2}$ from the manifold of such events $\mathcal{M}_{2}$, and a discrete, uniform ‘ring-like’ radiation pattern $\mathcal{U}_{N}$. The back-to-back configuration was first mentioned and computed by ATLAS in Ref. STDM-2020-20, but never in earlier phenomenological studies. Additionally, we consider several different distance metrics to illustrate how the choice of metric can affect the sensitivity of observables that we construct to distinct QCD physics processes.
@@ 2 Definition of event observables with EMD
This construction can be generalized to probe other global event geometries or specific structures by comparing the collider event to reference geometries. While there are many interesting candidates, we focus on two extremal examples: isotropic and dipole-like (back-to-back), as illustrated in Fig. 1. In the original work, the reference geometry was chosen to be a uniform (or, quasi-uniform, for computational tractability) distribution Cesarotti:2020hwb, but we will also consider the dipole geometry that was first shown in Ref. STDM-2020-20 to be an observable comparable to thrust. Additionally, we consider several distance measures to demonstrate how the sensitivity of these observables can be tuned to identify specific features of a geometry, e.g. by only considering transverse information in hadron collisions due to the unknown boost along the beamline. [^4] For the interested reader: Ref. STDM-2020-20 conducts some studies with longitudinal information included.
@@ 2.1 Reference Geometries
where $N$ is sufficiently large to approximate the infinite particle distribution, such that random rotations on the order of $\Delta\phi\sim 2\pi/N$ do not effect the distribution of the observable. For this work, we use $N=256$, such that $\mathcal{U}\approx\mathcal{U}_{256}$, shown in Fig. 2. We suppress these subscripts for notational simplicity, and will refer to the quasi-isotropic event as $\mathcal{U}$.
@@
where $\theta$ is the relative angle of the event, and the factors of $1/2$ are from normalizing each particle energy by the total energy. One additional aspect of determining the dipole distance $\mathcal{D}_{\mathcal{M}_{2}}$is to minimize the EMD over the orientation of the reference geometry with respect to the collider event. This is analogous—but not identical—to finding the thrust axis, or the axis along which the event is most collimated. In practice, the minimization over this angle is performed numerically using standard libraries when computing the manifold distance, e.g. Python Optimal Transport flamary2017pot, which is used throughout this result.
@@ 2.4 Analytics
Figure 5: Manifold distances as a function of NN-particle symmetric events (meaning equal energy, evenly distributed particles). The values of event isotropy are shown on the left and exhibit a monotonic behavior, all maximized for the 2-particle ($\mathcal{M}_{2}$) configuration, decreasing as the NN-particle event approaches the continuous limit. The values of (unnormalized) dipole distance $\mathcal{D}_{\mathcal{M}_{2}}$, shown on the right, are not monotonic, and jump around before asymptoting to the infinite particle limit. We show the unnormalized dipole distance values simply for ease of viewing. The dashed lines indicate the value of dipole distance computed for $N=\infty$. To the right of the solid line we only plot the value for NN divisible by 2 for ease of analytic computation.
The range of the manifold distances is closed and well defined for events with sets of massless particles and zero net momentum (‘recoil-corrected,’ following the discussion in Ref. Banfi:2010xy). Therefore, for ease of interpretation, subtract any unbalanced momentum from the event and normalize observables to their maximal value such that their domain is between 0 to 1. However, it is not immediately obvious what the maximally distant radiation pattern is. To show this, we plot the manifold distances as a function of symmetric $N$-particle radiation patterns $\mathcal{U}_{N}$ in Fig. 5
@@
The computation is less straightforward for dipole distance. For distance measures of $\beta\geq 1$, the maximally-distant configuration to the back-to-back reference geometry is the three-prong event. However, for $\beta=0.5$, we see in Fig. 5b that the maximally distant configuration from the dipole event is the uniform radiation pattern.
@@
$\bar{\mathcal{I}}\equiv 1$-$\mathcal{I}$ (11)
@@ 3.1 Monte Carlo setups and selections
The first sample consists of groups generated using the default Pythia group of hard QCD $2\rightarrow 2$ processes, which we call the ‘Hard QCD’ (‘HQCD’) sample. The implementation of the anti-$k_{t}$ algorithm with radius $R=0.4$ Cacciari:2008gp from FastJet is used Fastjet; Cacciari:2005hq to reconstruct jets from stable final-state particles with a transverse momentum above $500$ MeV and within $|y|<4.9$, excluding neutrinos. Such a requirement approximates the acceptance of detectors such as ATLAS or CMS at the LHC. The events in this sample are required to have at least two anti-$k_{t}$ jets with a $p_{T}$ above $60$ GeV and rapidity $|y|<4.5$ to be included in the study. Additional jets are also retained if they are present and satisfy the same criteria.
@@
In this figure, we note a few important features of the distributions. As expected, both observables demonstrate that the SQCD sample is more isotropic than the HQCD sample. As we have stated previously, manifold distances (and similar observables like thrust) have a greater dynamic range when the distance is computed with respect to a similar reference geometry than a very distant one: being ‘far’ from a radiation pattern is often a more degenerate mapping into a manifold distance. This illustrates how important the choice of reference geometry is for signal discrimination: a signal well-described by the underlying geometry will have support across the observable’s range, while a background poorly matched by the reference geometry will be pushed towards the extremal values, as indicated by the dashed lines at the mean of the distribution in Fig. 6. This feature reinforces the notion that ‘further’ events tend to saturate more quickly than ‘nearby’ events.
@@
[…]
@@
We start by looking at the HQCD samples in Fig. 8. It is immediately evident that for all values of $\beta$, the distributions are much more correlated for event isotropy than for dipole distance. We conjecture that this is due to the large underlying symmetry of the reference geometry for event isotropy. On average the distances traveled by particles to get to the reference geometry of a dipole rather than the isotropic distribution are greater, and the non-linearities in the EMD calculation are more evident.
@@
We also note the general trend that when the values of $\beta$ are more different, the observables become less directly correlated (i.e. there is more spread in the 2d distribution). This is a statement underscoring how different choices of $\beta$ fundamentally measure different features of the radiation pattern: high $\beta$ is insensitive to small movements, whereas smaller $\beta$ weights small and large distances increasingly equally. The increased spread demonstrates that, depending on the analysis at hand, studying two or more values of $\beta$ might lead to improved performance.
Figure 8: Plots of the 2d distributions of $\mathcal{D}_{\mathcal{M}_{2}}$ (top) and $\bar{\mathcal{I}}$ (bottom) for the HQCD sample. As a simple analytic model of the underlying behavior, we plot the curve for the maximally symmetric finite-multiplicity event ($\bar{\mathcal{I}}$  of $\mathcal{U}_{N}$) as a dashed line; this line represents the exact correlation we would expect if the samples were maximally symmetric. We see that $\bar{\mathcal{I}}$ is always more correlated, and that the correlations are less between high $\beta$ and low $\beta$.
@@
Now, let’s move our attention to the SQCD samples to confirm our statements are generic to the observables, and not specific to dijet-like radiation patterns: the same plots are shown for SQCD in Fig. 9. Again, we generally find that dipole distance is less correlated than event isotropy, except for the case of $\beta=0.5$ and $\beta=1$. The argument for this behavior is the same as that for the HQCD samples, but with the difference that for sufficiently small $\beta$, even little distances become relevant in the event isotropy calculation when the collider event is also somewhat isotropic. As we increase $\beta$, the distance weight for nearby particles becomes so small that the distributions become highly correlated. From the plots in Fig. 9, we see that the correlation collapses to the curve for the maximally isotropic value for a fixed number of particles.
Overall, we can conclude that the choice of $\beta$ when computing event isotropy doesn’t change the analysis as significantly as the choice of $\beta$ for dipole distance. The biggest exception we see to this statement is small $\beta$ for quasi-isotropic samples can produce different results as it is more sensitive to small perturbations from the reference geometry. Therefore, we recommend that unless there is a particular motivation to use a specific $\beta$, one should use whichever is least computationally expensive for event isotropy666i.e. avoiding square roots and other expensive operations when possible.. However, the choice of $\beta$, especially large vs. small, can noticeably change the distribution of dipole distance. The choice of $\beta$ should then be explored to optimize the analysis at hand.
@@
[…]
@@ 3.4 Particle and Jet Multiplicity
$f(\bar{\mathcal{I}},N)=\frac{\bar{\mathcal{I}}}{N}$ (12)
that is just the ratio of the event isotropy $\bar{\mathcal{I}}$ to the number of particles $N$ is a highly efficient measure. As shown in Fig. 12, we see that the spectra of the HQCD and SQCD sample are nearly completely separated, with an AUC value >> 0.99. We tested similar measures for the ratio and product of $\bar{\mathcal{I}}$ and $\mathcal{D}_{\mathcal{M}_{2}}$, but improvements to the AUC were negligible. A more complicated observable (e.g. one constructed with machine learning techniques hepmllivingreview) may still exist. Generally, when observables can probe different features of events, a new function with the observables as inputs of can serve as a highly efficient characterization measure Thaler:2010tr; Moult:2016cvt; Komiske:2017aww.
Figure 12: The HQCD and SQCD spectra of the ratio between event isotropy $\bar{\mathcal{I}}$ and particle number NN.
Alternatively, jet multiplicity reflects a more global picture of the radiation pattern as jets have a finite size radius. Thus an event with large jet multiplicity suggests that radiation has gone in several distinct directions, whereas a high-multiplicity event could indicate only that there are many collimated particles in a single jet. To study this correlation, we plot the distribution of the HQCD sample with $N_{j}\geq$ 2, 3, 5, and 7 in Fig. 13 along with the SQCD sample for reference. We see that jet multiplicity tracks the global features of the event more reliably, as the distribution in both event-shape observables shifts towards the SQCD sample, reinforcing one’s intuition that higher jet-multiplicity indicates a more isotropic event.